Manuel Kauers Hagenberg, May 12, 2003 RISC-Linz

Tenses II

Präteritum

Präteritum

• Def.: Let Xen be a regular German verb. Then the forms

Person Sg  Pl
1.  Xte Xten
2.  Xtest Xtet
3.  Xte Xten

define the Präteritum forms of Xen.

• Präteritum is the second past tense in German.

• Examples:

Er sagte während der ganzen Stunde kein Wort. Wir arbeiteten das ganze Wochenende.

• As with the past, many verbs are irregular. For instance,

Inf  ich  du  er  wir  ihr  sie
sein  war  warst  war  waren  wart  waren
haben hatte hattest hatte hatten hattet hatten
nehmen nahm nahmst nahm nahmen nahmt nahmen
gehen  ging gingst ging gingen gingt gingen
singen sang sangst sang sangen sangt sangen

• Examples:

Gestern war kein Deutschkurs. Am Wochenende besuchten wir ein Konzert.

• If A is the time of speaking, you may use Präteritum to formulate things that happend at time B before A.

You may use the Perfekt to formulate any event B before another event A, even when A is not the time of speaking.

Hence, Perfekt is “more universal” than Präteritum.

Exercise

• Convert the following sentences from Präteritum to Perfekt.

Frau Mayr kochte Kaffee. Ihre Enkelkinder besuchten sie. Früher wohnte Frau Mayr nicht in Hagenberg. Als sie nach Hagenberg kam, gab es das RISC noch nicht.

• Convert the following sentences from Perfekt to Präteritum.

Christian ist nach Insbruck gegangen. Er arbeitete vorher in der Theorema Gruppe und in der Gruppe von Josef. Er leitete den RISC-Deutschkurs. Einige Studenten lernten bei ihm Deutsch.

Irregularities

• The irregularity of most irregular verbs consists in a change of the vocal called Ablaut.

• What has to be learned (sorry) are the infinitive, the first person singular in Präteritum, and the PPP. All other forms can be built out of them.

• Examples: sprechen  sprach  gesprochen helfen  half  geholfen nehmen  nahm  genommen schwimmen schwamm geschwommen singen  sang  gesungen trinken  trank  getrunken gehen  ging  gegangen sehen  sah  gesehen lesen  las  gelesen essen  aß  gegessen

• The irregular verbs which belong to the Ablaut-class all have a PPP starting with ge- and ending in -en.

• There are sometimes additional differences besides the Ablaut, i.e., additional consonants are put between two vocals to make a form easier to pronounce.

• Of course, there are also some verbs which are completely irregular. (After all, we are dealing with German.)

• Examples:

sein  war  gewesen
haben  hatte  gehabt
vergessen vergaß  vergessen
entwickeln entwickelte entwickelt
bekommen bekam  bekommen
gewinnen gewann  gewonnen

Exercise

• Longterm Exercise: Make a list of all verbs you know, check if they are irregular, and collect the irregular forms in a table as above.